# Textual Knowledge Matters: Cross-Modality Co-Teaching

### v1

@@ 3.4 Cross-modality Co-teaching
In light of the categorical descriptive texts generated during the RTG phase, the primary challenge lies in effectively utilizing them to train an accurate GCD model. To address this, we develop two-stream parametric classifiers for visual and textual modalities. Recognizing that the intrinsic differences between the modalities naturally meet the requirements for the model disparity in co-teaching strategies Han et al. (2018), we propose a Cross-modal Co-Teaching (CCT) phase to realize the mutual benefits of the visual and textual classifiers. However, a notable challenge arises due to the lack of annotations during the co-teaching process: the classifiers for each modality are often misaligned. This misalignment implies that consistent class indexes across the two modalities cannot be guaranteed, posing a significant obstacle to effective co-teaching. We introduce two preliminary stages to overcome this issue: a warm-up stage and a class-aligning stage, designed to establish modality-specific classifiers while ensuring their alignment.
@@
$\mathcal{L}_{\text{base}}=\mathcal{L}^{\mathbf{I}}_{\text{base}}+\mathcal{L}^{\mathbf{T}}_{\text{base}},$ (3.4)
where both components are implemented by Eq. 3.3. The difference is that $\mathcal{L}^{\mathbf{T}}_{\text{base}}$ is calculated by the text model and textual description.
@@
$\mathcal{L}_{\text{con}}=\frac{1}{|B|}\sum_{i\in B}-\log\frac{\exp\left(\mathbf{f}_{i}^{\mathbf{I}}\cdot(\mathbf{f}_{i}^{\mathbf{T}})^{\top}/\tau_{c}\right)}{\sum_{j}\mathbb{I}_{[j\neq i]}\exp\left(\mathbf{f}_{i}^{\mathbf{I}}\cdot(\mathbf{f}_{j}^{\mathbf{T}})^{\top}/\tau_{c}\right)},$ (3.5)
@@
Stage I: Warm-up. The image and text models initially undergo a warm-up training stage with $e_{w}$ epochs, as depicted in Fig 4(a). During this stage, both the text model and image model are trained using the basic loss (Eq. 3.4) and cross-modal contrastive loss (Eq. 3.5). This warm-up training aims to enable both models to adequately learn category knowledge from their respective modality data, thereby developing modality-specific category perceptions.
Stage II: Class-aligning. Subsequently, to rectify the class misalignment between image and text classifiers, a class-aligning stage with $e_{a}$ epochs is introduced, as illustrated in Fig 4(b). This stage aims to align the classifiers of two modalities. To this end, we select high-confidence samples based on the text model and use them to guide the training of the image model. Specifically, for each category $k$, we select the top $s$ samples that have the highest probabilities in the $k$-th category produced by the text model, formulated as $\operatorname{Top}_{k}^{\mathbf{T}}(s)$. In addition, for each sample $i$ in $\operatorname{Top}_{k}^{\mathbf{T}}(s)$, we assign it with the hard pseudo label of $\widehat{y}_{i}^{\mathbf{T}}=k$. Thus, besides the basic and contrastive losses, we additionally train the image model by the pseudo-labeling loss:
$\mathcal{L}_{\text{p}}^{\mathbf{I}}=\frac{1}{|B^{s}|}\sum_{i\in B^{s}}\ell(\widehat{y}_{i}^{\mathbf{T}},\bm{p}_{i}^{\mathbf{I}}),$ (3.6)
@@
Stage III: Co-teaching. Following the warm-up and class-aligning stages, we establish cross-modality co-teaching to facilitate the mutual benefit between the image and text models, as shown in Fig 4(c). Similar to the class-aligning stage, we also generate high-confidence samples from the image model, which is defined as $\operatorname{Top}_{k}^{\mathbf{I}}(s)$. For each sample $i$ in $\operatorname{Top}_{k}^{\mathbf{I}}(s)$, the hard pseudo label is $\widehat{y}_{i}^{\mathbf{I}}=k$, and thus the pseudo-labeling loss for the text classifier is formulated as:
$\mathcal{L}_{\text{p}}^{\mathbf{T}}=\frac{1}{|B^{s}|}\sum_{i\in B^{s}}\ell(\widehat{y}_{i}^{\mathbf{I}},\bm{p}_{i}^{\mathbf{T}}).$ (3.7)
@@
$\mathcal{L}=\mathcal{L}_{\text{base}}+\mathcal{L}_{\text{con}}+\mathcal{L}_{\text{p}}^{\mathbf{I}}+\mathcal{L}_{\text{p}}^{\mathbf{T}}.$ (3.8)
@@
$\bm{P}_{i}=\bm{p}_{i}^{\mathbf{I}}+\bm{p}_{i}^{\mathbf{T}}.$ (3.9)
The predicted category is the one that achieves the largest value in $\bm{P}_{i}$, i.e., $\arg\max\bm{P}_{i}$. This integrated approach leverages the strengths of both modalities, leading to more accurate classification.
@@ 4 Experiments
Datasets. We conduct comprehensive experiments on four generic image classification datasets, i.e., CIFAR-10, CIFAR-100 Krizhevsky et al. (2009), ImageNet-100, and ImageNet-1K Deng et al. (2009), as well as on four fine-grained image classification datasets, i.e., CUB Wah et al. (2011), Stanford Cars Krause et al. (2013b), Oxford Pets Parkhi et al. (2012), and Flowers102 Nilsback & Zisserman (2008). Following Vaze et al. (2022), for each dataset, we select half of the classes as the known classes and the remaining as the unseen classes. We select 50% of samples from the known classes as the labeled dataset $\mathcal{D}_{L}$ and regard the remaining samples as unlabeled dataset $\mathcal{D}_{U}$ that are from both known and unseen classes. Refer to the Appendix Appendix material for details about the datasets.
Evaluation Protocol. Following Vaze et al. (2022), we use the clustering accuracy (ACC) to evaluate the performance of each algorithm. Specifically, the ACC is calculated on the unlabeled data, formulated as $ACC=\frac{1}{|\mathcal{D}_{U}|}\sum_{i=1}^{|\mathcal{D}_{U}|}\mathbb{I}(y_{i}=C(\overline{y}_{i}))$, where $\overline{y}=\arg\max\bm{P}$ represents the predicted labels and $y$ denotes the ground truth. The $C$ denotes the optimal permutation, aligning predicted cluster assignments with the actual class labels.
Implementation Details. We construct the tag lexicon by incorporating tags from semantic segmentation and object detection datasets Gupta et al. (2019); Lin et al. (2014); Kuznetsova et al. (2020), image classification datasets Russakovsky et al. (2015); Li et al. (2022a); Cimpoi et al. (2014); Parkhi et al. (2012); Bossard et al. (2014); Xiao et al. (2010); Krause et al. (2013a) and visual genome dataset Krishna et al. (2017). The CLIP Radford et al. (2021) model with a ViT-B-16 architecture serves as the backbone, ensuring a fair comparison to methods utilizing the DINO Caron et al. (2021) model with the same architecture. A linear classifier layer is added after the backbone for each modality. During training, only the last layers of both the text and image encoders in the backbone are fine-tuned, using a learning rate of 0.0005. The learning rates for classifiers start at 0.1 and decrease following a cosine annealing schedule. We train the model for 200 epochs on all datasets, using a batch size of 128 and processing dual views of randomly augmented images and texts. The temperature parameters $\tau_{s}$ and $\tau_{u}$ are set at 0.1 and 0.05, respectively. $\tau_{t}$ starts at 0.035 and decreases to 0.02, except for CIFAR-10 where it varies from 0.07 to 0.04 due to fewer classes. We use the logit scale value to set $\tau_{c}$ as in the CLIP model. The balancing factor $\lambda$ is set to 0.2. Warm-up and class-aligning stages last for 10 and 5 epochs, respectively, denoted as $e_{w}=10$ and $e_{a}=5$. For selecting high-confidence samples, we choose $s=r*\frac{|\mathcal{D}_{U}|}{K}$ samples per category, with $r$ set at 0.6. In line with SimGCD Wen et al. (2023), we conduct each experiment three times, and the average result over the three runs is reported.
@@ 4.1 Comparison with the State of the Art
We conduct a comparative analysis between our proposed TextGCD and leading methods in GCD, as delineated in Tab. 1 and Tab. 2. These methods encompass GCD Vaze et al. (2022), SimGCD Wen et al. (2023), DCCL Pu et al. (2023), GPC Zhao et al. (2023), PromptCAL Zhang et al. (2023), CLIP-GCD Ouldnoughi et al. (2023) and SPTNet Wang et al. (2023). Except for CLIP-GCD Ouldnoughi et al. (2023), which uses CLIP as the backbone, the others utilize the DINO as the backbone. To better ensure the fairness of our comparison, we also reproduce SimGCD by using CLIP as the backbone. Clearly, our TextGCD outperforms all compared methods in terms of All accuracy on all datasets. Importantly, compared to the best competitor, SimGCD, our method achieves a significant improvement on the challenging ImageNet-1K and all the fine-grained datasets. Specifically, our method outperforms SimGCD by 7.7% in All accuracy and 11.4% in New accuracy on the ImageNet-1K dataset, and, by 11.1% in All accuracy and 12.7% in New accuracy averaged on fine-grained datasets. This demonstrates the superiority of our method over existing visual-based methods. In comparison to CLIP-GCD, which merely concatenates features from both modalities, TextGCD achieves higher results. For instance, TextGCD surpasses CLIP-GCD by 13.8%, 16.3%, and 10.9% in terms of All accuracy on the CUB, Stanford Cars, and Flowers102 datasets, respectively.
@@
GCD Vaze et al. (2022) DINO 91.5 97.9 88.2 73.0 76.2 66.5 74.1 89.8 66.3 52.5 72.5 42.2
SimGCD Wen et al. (2023) DINO 97.1 95.1 98.1 80.1 81.2 77.8 83.0 93.1 77.9 57.1 77.3 46.9
DCCL Pu et al. (2023) DINO 96.3 96.5 96.9 75.3 76.8 70.2 80.5 90.5 76.2 - - -
GPC Zhao et al. (2023) DINO 92.2 98.2 89.1 77.9 85.0 63.0 76.9 94.3 71.0 - - -
PromptCAL Zhang et al. (2023) DINO 97.9 96.6 98.5 81.2 84.2 75.3 83.1 92.7 78.3 - - -
SPTNet Wang et al. (2023) DINO 97.3 95.0 98.6 81.3 84.3 75.6 85.4 93.2 81.4 - - -
CLIP-GCD Ouldnoughi et al. (2023) CLIP 96.6 97.2 96.4 85.2 85.0 85.6 84.0 95.5 78.2 - - -
SimGCD Wen et al. (2023) CLIP 96.6 94.7 97.5 81.6 82.6 79.5 86.1 94.5 81.9 48.2 72.7 36.0
@@
GCD Vaze et al. (2022) DINO 51.3 56.6 48.7 39.0 57.6 29.9 80.2 85.1 77.6 74.4 74.9 74.1
SimGCD Wen et al. (2023) DINO 60.3 65.6 57.7 53.8 71.9 45.0 87.7 85.9 88.6 71.3 80.9 66.5
DCCL Pu et al. (2023) DINO 63.5 60.8 64.9 43.1 55.7 36.2 88.1 88.2 88.0 - - -
SPTNet Wang et al. (2023) DINO 65.8 68.8 65.1 59.0 79.2 49.3 - - - - - -
CLIP-GCD Ouldnoughi et al. (2023) CLIP 62.8 77.1 55.7 70.6 88.2 62.2 - - - 76.3 88.6 70.2
SimGCD Wen et al. (2023) CLIP 62.0 76.8 54.6 75.9 81.4 73.1 88.6 75.2 95.7 75.3 87.8 69.0
@@ 4.2 Ablation Study
Components Ablation. In Tab. 3, we present an ablation study of the key components in TextGCD, specifically cross-modality co-teaching, contrastive learning, and soft voting. Beginning with the Baseline method, we incrementally incorporate these three components into the framework to evaluate their impact on the performance of both image and text classifiers. We also report the results of 1) employing the first tag derived from the visual lexicon as the GCD labels and 2) performing $k$-means MacQueen et al. (1967) using the image or text features without training. It becomes evident that the “First-Tag” and “$k$-means” methods yield inferior results compared to the baseline. On the other hand, the “$k$-means” method and the baseline results demonstrate that using the textual information achieves clearly higher results than using the image cues, thereby highlighting the superior quality of the generated textual descriptions. Compared to the baseline, integrating our proposed co-teaching approach markedly elevates the performance across both modalities, particularly notable within the context of the fine-grained CUB dataset. This substantiates the efficacy of our co-teaching strategy in fostering synergistic learning and collaborative advancement between text and image models. Furthermore, image-text contrastive learning consistently increases the All and New accuracies, underlining the necessity of aligning the two modalities. Ultimately, soft voting can further improve performance, illustrating that the integration of information from both modalities leads to more precise category discrimination.
@@ 4.3 Evaluation
Auxiliary Model Evaluation. In Tab. 8, we explore the effect of substituting the auxiliary model with FLIP Li et al. (2023b) and CoCa Yu et al. (2022). Results show that our method consistently produces high performance on both datasets and that using a more powerful model (FLIP) leads to higher results.
@@
CLIP Radford et al. (2021) ViT-H-14 85.7 86.3 84.6 76.6 80.6 74.7
CoCa Yu et al. (2022) ViT-L-14 85.2 85.8 83.8 73.6 81.7 69.5
FLIP Li et al. (2023b) ViT-G-14 87.6 87.7 87.5 79.5 83.5 77.4

### v2

@@ 3.4 Cross-modality Co-teaching
In light of the categorical descriptive texts generated during the RTG phase, the primary challenge lies in effectively utilizing them to train an accurate GCD model. To address this, we develop two-stream parametric classifiers for visual and textual modalities. Recognizing that the intrinsic differences between the modalities naturally meet the requirements for the model disparity in co-teaching strategies [12], we propose a Cross-modal Co-Teaching (CCT) phase to realize the mutual benefits of the visual and textual classifiers. However, a notable challenge arises due to the lack of annotations during the co-teaching process: the classifiers for each modality are often misaligned. This misalignment implies that consistent class indexes across the two modalities cannot be guaranteed, posing a significant obstacle to effective co-teaching. We introduce two preliminary stages to overcome this issue: a warm-up stage and a class-aligning stage, designed to establish modality-specific classifiers while ensuring their alignment.
@@
$\mathcal{L}_{\text{base}}=\mathcal{L}^{\mathbf{I}}_{\text{base}}+\mathcal{L}^{\mathbf{T}}_{\text{base}},$ (4)
where both components are implemented by Eq. 3. The difference is that $\mathcal{L}^{\mathbf{T}}_{\text{base}}$ is calculated by the text model and textual description.
@@
$\mathcal{L}_{\text{con}}=\frac{1}{|B|}\sum_{i\in B}-\log\frac{\exp\left(\mathbf{f}_{i}^{\mathbf{I}}\cdot(\mathbf{f}_{i}^{\mathbf{T}})^{\top}/\tau_{c}\right)}{\sum_{j}\mathbb{I}_{[j\neq i]}\exp\left(\mathbf{f}_{i}^{\mathbf{I}}\cdot(\mathbf{f}_{j}^{\mathbf{T}})^{\top}/\tau_{c}\right)},$ (5)
@@
Stage I: Warm-up. The image and text models initially undergo a warm-up training stage with $e_{w}$ epochs, as depicted in Fig 4(a). During this stage, both the text model and image model are trained using the basic loss (Eq. 4) and cross-modal contrastive loss (Eq. 5). This warm-up training aims to enable both models to adequately learn category knowledge from their respective modality data, thereby developing modality-specific category perceptions.
@@
Stage II: Class-aligning. Subsequently, to rectify the class misalignment between image and text classifiers, a class-aligning stage with $e_{a}$ epochs is introduced, as illustrated in Fig 4(b). This stage aims to align the classifiers of two modalities. To this end, we select high-confidence samples based on the text model and use them to guide the training of the image model. Specifically, for each category $k$, we select the top $s$ samples that have the highest probabilities in the $k$-th category produced by the text model, formulated as $\operatorname{Top}_{k}^{\mathbf{T}}(s)$. In addition, for each sample $i$ in $\operatorname{Top}_{k}^{\mathbf{T}}(s)$, we assign it with the hard pseudo label of $\hat{y}_{i}^{\mathbf{T}}=k$. Thus, besides the basic and contrastive losses, we additionally train the image model by the pseudo-labeling loss:
$\mathcal{L}_{\text{p}}^{\mathbf{I}}=\frac{1}{|B^{s}|}\sum_{i\in B^{s}}\ell(\hat{y}_{i}^{\mathbf{T}},\boldsymbol{p}_{i}^{\mathbf{I}}),$ (6)
@@
Stage III: Co-teaching. Following the warm-up and class-aligning stages, we establish cross-modality co-teaching to facilitate the mutual benefit between the image and text models, as shown in Fig 4(c). Similar to the class-aligning stage, we also generate high-confidence samples from the image model, which is defined as $\operatorname{Top}_{k}^{\mathbf{I}}(s)$. For each sample $i$ in $\operatorname{Top}_{k}^{\mathbf{I}}(s)$, the hard pseudo label is $\hat{y}_{i}^{\mathbf{I}}=k$, and thus the pseudo-labeling loss for the text classifier is formulated as:
$\mathcal{L}_{\text{p}}^{\mathbf{T}}=\frac{1}{|B^{s}|}\sum_{i\in B^{s}}\ell(\hat{y}_{i}^{\mathbf{I}},\boldsymbol{p}_{i}^{\mathbf{T}}).$ (7)
@@
$\mathcal{L}=\mathcal{L}_{\text{base}}+\mathcal{L}_{\text{con}}+\mathcal{L}_{\text{p}}^{\mathbf{I}}+\mathcal{L}_{\text{p}}^{\mathbf{T}}.$ (8)
@@
$\boldsymbol{P}_{i}=\boldsymbol{p}_{i}^{\mathbf{I}}+\boldsymbol{p}_{i}^{\mathbf{T}}.$ (9)
The predicted category is the one that achieves the largest value in $\boldsymbol{P}_{i}$, i.e., $\arg\max\boldsymbol{P}_{i}$. This integrated approach leverages the strengths of both modalities, leading to more accurate classification.
@@ 4 Experiments
Datasets. We conduct comprehensive experiments on four generic image classification datasets, i.e., CIFAR-10, CIFAR-100 [18], ImageNet-100, and ImageNet-1K [8], as well as on four fine-grained image classification datasets, i.e., CUB [41], Stanford Cars [15], Oxford Pets [31], and Flowers102 [28]. Following [40], for each dataset, we select half of the classes as the known classes and the remaining as the unseen classes. We select 50% of samples from the known classes as the labeled dataset $\mathcal{D}_{L}$ and regard the remaining samples as unlabeled dataset $\mathcal{D}_{U}$ that are from both known and unseen classes.
Evaluation Protocol. Following [40], we use the clustering accuracy (ACC) to evaluate the performance of each algorithm. Specifically, the ACC is calculated on the unlabeled data, formulated as $ACC=\frac{1}{|\mathcal{D}_{U}|}\sum_{i=1}^{|\mathcal{D}_{U}|}\mathbb{I}(y_{i}=C(\overline{y}_{i}))$, where $\overline{y}=\arg\max\boldsymbol{P}$ represents the predicted labels and $y$ denotes the ground truth. The $C$ denotes the optimal permutation.
Implementation Details. We construct the tag lexicon by incorporating tags from semantic segmentation and object detection datasets [11, 24, 19], image classification datasets [37, 20, 7, 31, 4, 44, 16] and visual genome dataset [17]. The CLIP [33] model with a ViT-B-16 architecture serves as the backbone, ensuring a fair comparison to methods utilizing the DINO [6] model with the same architecture. A linear classifier layer is added after the backbone for each modality. During training, only the last layers of both the text and image encoders in the backbone are fine-tuned, using a learning rate of 0.0005. The learning rates for classifiers start at 0.1 and decrease following a cosine annealing schedule. We train the model for 200 epochs on all datasets, using a batch size of 128 and processing dual views of randomly augmented images and texts. The temperature parameters $\tau_{s}$ and $\tau_{u}$ are set at 0.1 and 0.05, respectively. $\tau_{t}$ starts at 0.035 and decreases to 0.02, except for CIFAR-10 where it varies from 0.07 to 0.04 due to fewer classes. We use the logit scale value to set $\tau_{c}$ as in the CLIP model. The balancing factor $\lambda$ is set to 0.2. Warm-up and class-aligning stages last for 10 and 5 epochs, respectively, denoted as $e_{w}=10$ and $e_{a}=5$. For selecting high-confidence samples, we choose $s=r*\frac{|\mathcal{D}_{U}|}{K}$ samples per category, with $r$ set at 0.6. The average result over the three runs is reported.
@@ 4.1 Comparison with the State of the Art
We conduct a comparative analysis between our proposed TextGCD and leading methods in GCD, as delineated in Tab. 1 and Tab. 2. These methods encompass GCD [40], SimGCD [43], DCCL [32], GPC [51], PromptCAL [49], CLIP-GCD [30] and SPTNet [42]. Except for CLIP-GCD [30], which uses CLIP as the backbone, the others utilize the DINO as the backbone. To better ensure the fairness of our comparison, we also reproduce SimGCD by using CLIP as the backbone. Clearly, our TextGCD outperforms all compared methods in terms of All accuracy on all datasets. Importantly, compared to the best competitor, SimGCD, our method achieves a significant improvement on the ImageNet-1K and all the fine-grained datasets. Specifically, our method outperforms SimGCD by 7.7% in All accuracy and 11.4% in New accuracy on the ImageNet-1K dataset, and, by 11.1% in All accuracy and 12.7% in New accuracy averaged on fine-grained datasets. This demonstrates the superiority of our method over existing visual-based methods. In comparison to CLIP-GCD, which merely concatenates features from both modalities, TextGCD achieves higher results. For instance, TextGCD surpasses CLIP-GCD by 13.8%, 16.3%, and 10.9% in terms of All accuracy on the CUB, Stanford Cars, and Flowers102 datasets, respectively.
@@
GCD DINO 91.5 97.9 88.2 73.0 76.2 66.5 74.1 89.8 66.3 52.5 72.5 42.2
SimGCD DINO 97.1 95.1 98.1 80.1 81.2 77.8 83.0 93.1 77.9 57.1 77.3 46.9
DCCL DINO 96.3 96.5 96.9 75.3 76.8 70.2 80.5 90.5 76.2 - - -
GPC DINO 92.2 98.2 89.1 77.9 85.0 63.0 76.9 94.3 71.0 - - -
PromptCAL DINO 97.9 96.6 98.5 81.2 84.2 75.3 83.1 92.7 78.3 - - -
SPTNet DINO 97.3 95.0 98.6 81.3 84.3 75.6 85.4 93.2 81.4 - - -
CLIP-GCD CLIP 96.6 97.2 96.4 85.2 85.0 85.6 84.0 95.5 78.2 - - -
SimGCD CLIP 96.6 94.7 97.5 81.6 82.6 79.5 86.1 94.5 81.9 48.2 72.7 36.0
@@
GCD DINO 51.3 56.6 48.7 39.0 57.6 29.9 80.2 85.1 77.6 74.4 74.9 74.1
SimGCD DINO 60.3 65.6 57.7 53.8 71.9 45.0 87.7 85.9 88.6 71.3 80.9 66.5
DCCL DINO 63.5 60.8 64.9 43.1 55.7 36.2 88.1 88.2 88.0 - - -
SPTNet DINO 65.8 68.8 65.1 59.0 79.2 49.3 - - - - - -
CLIP-GCD CLIP 62.8 77.1 55.7 70.6 88.2 62.2 - - - 76.3 88.6 70.2
SimGCD CLIP 62.0 76.8 54.6 75.9 81.4 73.1 88.6 75.2 95.7 75.3 87.8 69.0
@@ 4.2 Ablation Study
Components Ablation. In Tab. 3, we present an ablation study of the key components in TextGCD, specifically cross-modality co-teaching, contrastive learning, and soft voting. Beginning with the Baseline method, we incrementally incorporate these three components into the framework to evaluate their impact on the performance of both image and text classifiers. We also report the results of 1) employing the first tag derived from the visual lexicon as the GCD labels and 2) performing $k$-means [25] using the image or text features without training. It becomes evident that the “First-Tag” and “$k$-means” methods yield inferior results compared to the baseline. On the other hand, the “$k$-means” method and the baseline results demonstrate that using the textual information achieves clearly higher results than using the image cues, thereby highlighting the superior quality of the generated textual descriptions. Compared to the baseline, integrating our proposed co-teaching approach markedly elevates the performance across both modalities, particularly notable within the context of the fine-grained CUB dataset. This substantiates the efficacy of our co-teaching strategy in fostering synergistic learning and collaborative advancement between text and image models. Furthermore, image-text contrastive learning consistently increases the All and New accuracies, underlining the necessity of aligning the two modalities. Soft voting can further improve performance, illustrating that the integration of information from both modalities leads to more precise category discrimination.
@@ 4.3 Evaluation
Auxiliary Model Evaluation. In Tab. 8, we explore the effect of substituting the auxiliary model with FLIP [23] and CoCa [47]. Results show that our method consistently produces high performance on both datasets and that using a more powerful model (FLIP) leads to higher results.
@@
CLIP ViT-H-14 85.7 86.3 84.6 76.6 80.6 74.7
CoCa ViT-L-14 85.2 85.8 83.8 73.6 81.7 69.5
FLIP ViT-G-14 87.6 87.7 87.5 79.5 83.5 77.4